$SPAD/src/input scope.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool scope.output
)set message test on
)set message auto off
)clear all
--S 1 of 3
showbug1():Void ==
 for i in 1..1 repeat
  z:="I'm OK"
  print(z)
  showbug2()
  print(z)
--R
--R   Function declaration showbug1 : () -> Void has been added to
--R      workspace.
--R                                                                   Type: Void
--E 1
--S 2 of 3
showbug2():Void ==
 for i in 1..1 repeat
  z:="I'm nasty"
--R
--R   Function declaration showbug2 : () -> Void has been added to
--R      workspace.
--R                                                                   Type: Void
--E 2
-- used to print:
-- I'm OK
-- I'm nasty
--S 3 of 3
showbug1()
--R
--R   Compiling function showbug2 with type () -> Void
--R   Compiling function showbug1 with type () -> Void
--R   "I'm OK"
--R   "I'm OK"
--R                                                                   Type: Void
--E 3

)spool )lisp (bye) ———-

4

References^{[1] nothing}

5